Solve the following loarithmic equations
log(x-5)=2
5+log2x=6

Answer by Alan3354(21583)

Solve the following loarithmic equations
log(x-5)=2
x-5 = 10^2 = 100
x = 105
-------------
5+log2x=6
log(2x) = 1
2x = 10
x = 5